OpenStudy (anonymous):

Evaluate 6+4y-x^2 for x=-5;y=6

OpenStudy (anonymous):

just plug in the x=-5 and y=6 use calculator.... 6+4(6)-(-5)^2=?

OpenStudy (anonymous):

that is 6+24-25=?

OpenStudy (anonymous):

did u get it? 24-25=-1... 6-1=5

OpenStudy (anonymous):

Wouldn't it be 6 + 24 +25 though since a negative and a negative makes a positive?

OpenStudy (anonymous):

no, that should be 6+24 - (+25)= 6+24-25=5

OpenStudy (anonymous):

there is a hiearchy on this, first is power, then multiplication or division then add or subtract

OpenStudy (anonymous):

Ok.Thanks.

OpenStudy (anonymous):

wc